OpenStudy (anonymous):

Identify the solution(s) of the system of equations, if any. 3x-2y=6 5x=5-25y

OpenStudy (anonymous):

well this is a tough u have to multiply both equations soo u can elemenate one x or y

OpenStudy (anonymous):

soo lets choose to eliminate x equation 1 * -5 : -15x+10y=34 equation 2 * 3 : 15x +75y =15

OpenStudy (anonymous):

add them together we have 85y=49 y=49/85

OpenStudy (anonymous):

something is wrong

OpenStudy (anonymous):

wouldn't equation 1 be -10y ?

OpenStudy (anonymous):

equation 1 -15x+10y=-34

OpenStudy (anonymous):

oh because its a negative and negative =+

OpenStudy (texaschic101):

5x = 5 - 25y x = 1 - 5y now sub 1 - 5y in for x in the other equation

OpenStudy (texaschic101):

3x - 2y = 6 3(1 - 5y) - 2y = 6 can you take it from here and solve for y ?

OpenStudy (texaschic101):

or you can continue doing elimination. I personally think substitution is sometimes easier

OpenStudy (anonymous):

equation 1 * -5 : -15x+10y=-34 equation 2 * 3 : 15x +75y =15 85y=-34+15

OpenStudy (anonymous):

it is 30 not 34 sorry

OpenStudy (anonymous):

Its ok- I didn't catch it either equation 1 * -5 : -15x+10y=-30 equation 2 * 3 : 15x +75y =15

OpenStudy (anonymous):

85y=-15, right?

OpenStudy (anonymous):

yup soo 85y=-15 y=-3/17

OpenStudy (anonymous):

by substitution 5x=5-25y 5x=5-(25*-3/17) x= 32/17

OpenStudy (anonymous):

gotcha! I appreciate how you explain things, very easy to follow!

OpenStudy (anonymous):

to confirm the answer we can use the other equation 3x-2y=6 and sub x=32/17 and y =-3/17 3*(32/17)-2(-3/17) we would find it equals 6

OpenStudy (anonymous):

glad i can help out :)